OpenStudy (anonymous):

Let T: R^2-> R^2 be the linear transformation that projects points onto the line y= 3x. Write down all the eigenvalues and eigenvectors of T. n explain how do u get them?

OpenStudy (jamesj):

Well intuitively (1,3) is an eigenvector. But to get at this formally, write down the matrix of T. \[T : i \to i + 3j\] and \[T : j \to \frac{1}{3} i + j\] Hence \[T = \left[\begin{matrix}1 & 1/3 \\ 3 & 1 \end{matrix}\right]\] Now solve for eigenvalues and -vectors as usual \[\det(A - \lambda I) = 0 \implies ...\]

OpenStudy (anonymous):

n again thank you

OpenStudy (jamesj):

The other vector that is intuitively an eigenvector is the 1-d basis of the subspace that gets sent to the null vector; that is every vector lying along the line perpendicular to y = 3x; that is, the line y = -x/3. Hence the basis for that subspace is (1,-1/3)

OpenStudy (jamesj):

So hopefully something like that comes out of the formal calculation. (and hopefully I haven't made a sign error!)